OpenStudy (anonymous):

can anyone help with simplifying the following expression? 1+ square root of 2 over 1 - square root of 2

OpenStudy (anonymous):

\[1+\sqrt[2]{1}-\sqrt{2}=2-\sqrt{2}\]

OpenStudy (anonymous):

no, let me write it again

OpenStudy (anonymous):

\[1 + \sqrt{2} \over 1-\sqrt{2}\]